# Almost equivalences between Tamarkin category and Novikov sheaves

### v2

@@ 5.2 Derived completeness
We will show the space $\operatorname{Hom}_{\mu^{\mathbb{R}}(*)}(\mathcal{E},\mathcal{F})$ for $\mathcal{E},\mathcal{F}$ is derived complete. The other cases will follow from similar arguments. In other words, it is enough to show the homotopy limit of the sequence
[…]
@@
By using the internal hom defined in [Tam18, Kuwal], we have
$\begin{split}\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\operatorname{Hom}_{\mu^{\mathbb{R}}(*)}(\mathcal{E},\mathcal{F})&=\operatorname{Hom}_{\mu^{\mathbb{R}}(*)}(\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c},\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathop{{\mathcal{H}}om}\nolimits^{\star_{\mathbb{R}}}(\mathcal{E},\mathcal{F}))\\ &=\operatorname{Hom}_{\mu^{\mathbb{O}}(*)}(\mathbb{K}_{t\geq 0},\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathop{{\mathcal{H}}om}\nolimits^{\star_{\mathbb{R}}}(\mathcal{E},\mathcal{F}))\\ \end{split}$ (5.12)
@@
[…]
@@
$\begin{split}\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathcal{G})&\cong\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathcal{G})\\ &\cong\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\mathcal{G})\\ &\cong\lim_{\begin{subarray}{c}\longleftarrow\\ c\rightarrow\infty\end{subarray}}\Gamma(\mathbb{K}_{\left\{t\geq c\right\}},\mathcal{G})\\ &\cong\Gamma(\lim_{\begin{subarray}{c}\longrightarrow\\ c\rightarrow\infty\end{subarray}}\mathbb{K}_{\left\{t\geq c\right\}},\mathcal{G})\cong 0.\end{split}$ (5.14)
This completes the proof. ∎
@@ 5.3 Morita functor
We first show the conservativity of the functor: If an object $\mu^{\mathbb{G}}(T^{*}\{*\})$ satisfies $\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},\mathcal{E})=0$, then $\mathcal{E}=0$. Actually, if $\mathcal{E}$ satisfies $\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},\mathcal{E})=0$, it means the vanishing of positive microsupport, which means $\mathcal{E}=0$.
@@
$\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\operatorname{Hom}(\mathbb{K},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\operatorname{Hom}(\mathbb{K}_{(-\infty,0)},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow.$ (5.19)
As in the proof of Lemma 5.2, we can see that $H^{1}(\operatorname{Hom}(\mathbb{K},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c}))\simeq 0$. Then one can go to the cohomology exact sequence as
$0\rightarrow H^{0}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\widehat{\bigoplus}_{c\in(-\infty,+\infty)}\mathbb{K}\rightarrow\widehat{\bigoplus}_{c\in(-\infty,0)}\mathbb{K}\rightarrow H^{1}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow 0$ (5.20)
Here the completion of ^$\widehat{\bigoplus}_{c\in(-\infty,0)}\mathbb{K}$ is taken in the direction to $\infty$- ∞ and 00 and the completion of ^$\widehat{\bigoplus}_{c\in(-\infty,0)}\mathbb{K}$ is taken in the direction to $\infty$- ∞ and +$\infty$+ ∞. Hence we conclude that
$H^{i}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\cong\begin{cases}&\Lambda_{0}\text{ if $i=0$}\\ &\operatorname{coker}\left(\bigoplus_{[1,0)}\mathbb{K}\rightarrow\widehat{\bigoplus}_{[1,0)}\mathbb{K}\right)\text{ if $i=1$}\\ &0\text{ otherwise}.\end{cases}$ (5.21)
where the morphism in the second line is a natural one. Then the almostization kill the degree one morphisms.
@@
$\operatorname{Hom}(1_{\mu},\bigoplus_{i\in I}T_{d_{i}}1_{\mu})\cong\widehat{\bigoplus}_{i\in I}\operatorname{Hom}(T_{c}1_{\mu},T_{d_{i}}1_{\mu}).$ (5.22)
@@
By using the standard argument (cf. [SS03]) and Step 2, for any object $\mathcal{E}\in\mu^{\mathbb{G}}(*)$, we can construct an exact triangle
$A\rightarrow\mathcal{E}\rightarrow B\xrightarrow{[1]}$ (5.23)
where $A$ is in the colimits of $\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu}$ and $B$ is in the right orthogonal of such colimits. In Step 1, we check that such an orthogonal is zero. Hence any $\mathcal{E}$ is in the colimits of $\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu}$. Since $\mathfrak{A}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong_{\mathrm{a}}L_{0}^{\mathbb{G}}$ by Lemma 5.3 and
$\operatorname{End}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong_{\mathrm{a}}\operatorname{End}(L_{0}^{\mathbb{G}})=L_{0}^{\mathbb{G}},$ (5.24)
we conclude that the functor $\mathfrak{A}$ is almost fully faithful.
Also, since $L_{0}^{\mathbb{G}}$ is a generator of $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ and $\mathfrak{A}$ is cocontinuous, the almost essential surjectivity follows. This completes the proof.
@@ 6.1 Reminders on the Lurie tensor product
We follow Volpe’s exposition [Vol21]. Let $Cat_{comp}$ be the category of the cocomplete categories and the morphisms are cocontinuous functors. For cocomplete categories $\mathcal{C},\mathcal{D}$, there exists a cocomplete category $\mathcal{C}\otimes_{L}\mathcal{D}$ with a functor $\mathcal{C}\times\mathcal{D}\rightarrow\mathcal{C}\otimes_{L}\mathcal{D}$ satisfying
$Fun(\mathcal{C}\times\mathcal{D},\mathcal{E})\cong Fun(\mathcal{C}\otimes_{L}\mathcal{D},\mathcal{E})$ (6.1)
where the left hand side denotes the functors preserving variable-wise colimits and the right hand side denotes the functors preserving colimits. The resulting category $\mathcal{C}\otimes_{L}\mathcal{D}$ is called Lurie’s tensor product. In the following, we simply denote $\otimes_{L}$ by $\otimes$.
@@
Lemma 6.1 ([Vol21, Corollary 2.30]).
Let $M$ and $N$ be manifolds.
@@
    We have an equivalence. $\mathrm{Sh}(M,\mathbb{K})\otimes\mathrm{Sh}(N,\mathbb{K})\cong\mathrm{Sh}(M\times N,\mathbb{K})$.
@@
    Let $\mathcal{C}$ be a presentable $\mathbb{K}$-linear category. Then we have $\mathrm{Sh}(M,\mathbb{K})\otimes\mathcal{C}\cong\mathrm{Sh}(M,\mathcal{C})$.

### v3

@@ 5.2 Derived completeness
We first show the space $\operatorname{Hom}_{\mu^{\mathbb{R}}(*)}(\mathcal{E},\mathcal{F})$ for $\mathcal{E},\mathcal{F}$ is derived complete. In other words, it is enough to show the homotopy limit of the sequence
$\cdots\xrightarrow{T}\operatorname{Hom}(\mathcal{E},\mathcal{F})\xrightarrow{T}\operatorname{Hom}(\mathcal{E},\mathcal{F})\xrightarrow{T}\operatorname{Hom}(\mathcal{E},\mathcal{F}).$ (5.10)
@@
By using the internal hom defined in [Tam18, Kuw24], we have
[…]
where we use Lemma 5.1(4) for the second equality.
@@
$\cdots\xrightarrow{T}\mathcal{G}\xrightarrow{T}\mathcal{G}\xrightarrow{T}\mathcal{G}.$ (5.13)
@@
$\begin{split}\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\mathcal{G})&\cong\lim_{\begin{subarray}{c}\longleftarrow\\ i\rightarrow\infty\end{subarray}}\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\mathcal{G})\\ &\cong\lim_{\begin{subarray}{c}\longleftarrow\\ c\rightarrow\infty\end{subarray}}\Gamma(\mathbb{K}_{\left\{t\geq c\right\}},\mathcal{G})\\ &\cong\Gamma(\lim_{\begin{subarray}{c}\longrightarrow\\ c\rightarrow\infty\end{subarray}}\mathbb{K}_{\left\{t\geq c\right\}},\mathcal{G})\cong 0.\end{split}$ (5.14)
Here, in the second equality, we use the following commutative diagram to transfer the $T$-action from $\mathcal{G}$ to $\mathbb{K}_{t\geq c}$:
${\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\mathcal{G})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T\circ(-)}\scriptstyle{\cong}$${\Gamma(\mathbb{K}_{\left\{t\geq 0\right\}},\mathcal{G})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\Gamma(\mathbb{K}_{\left\{t\geq n\right\}},\mathcal{G})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Gamma(\mathbb{K}_{\left\{t\geq n+1\right\}},\mathcal{G})}$ (5.15)
where the lower horizontal arrow is induced by the canonical morphism $\mathbb{K}_{t\geq n}\to\mathbb{K}_{t\geq n+1}$, and vertical arrows are induced by translations. This completes the proof.
For general $\mathbb{G}$, we can equip the $\mathbb{G}$-equivariance $\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E}$ (which is the left adjoint of the forgetful functor $\mu^{\mathbb{R}}(*)\to\mu^{\mathbb{G}}(*)$. See Lemma  5.1 (4) for a related assertion). This structure induces $\Lambda_{0}$-action on $\operatorname{Hom}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F})$, which is nothing but the structure of $\mathfrak{f}(\operatorname{Hom}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F}))$. Hence the derived completeness of $\operatorname{Hom}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F})$ is again the vanishing of the homotopy limit of the sequence
$\cdots\xrightarrow{T}\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F})\xrightarrow{T}\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F})\xrightarrow{T}\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}\mathcal{E},\mathcal{F}),$ (5.16)
which can be shown by the same argument as above. ∎
@@ 5.3 Morita functor
We first show the conservativity of the functor: If an object $\mathcal{E}\in\mu^{\mathbb{G}}(\{*\})$ satisfies $\operatorname{Hom}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},\mathcal{E})=0$, then $\mathcal{E}=0$. Suppose $\mathcal{E}$ satisfies $\operatorname{Hom}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},\mathcal{E})=0$. Since $\operatorname{Cone}(T^{c^{\prime}}\colon 1_{\mu}\to 1_{\mu})=\bigoplus_{c\in\mathbb{G}}\mathbb{K}_{c\leq t<c+c^{\prime}}$ for any $c^{\prime}>0$, we have
$\begin{split}0&=\operatorname{Hom}_{\mu^{\mathbb{G}}(*)}(\bigoplus_{c^{\prime\prime}\in\mathbb{R}/\mathbb{G}}T_{c^{\prime\prime}}\bigoplus_{c\in\mathbb{G}}\mathbb{K}_{c\leq t<c+c^{\prime}},\mathcal{E})\\ &\cong\operatorname{Hom}_{\mathrm{Sh}(\mathbb{R}_{t})}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{\widetilde{c}}\mathbb{K}_{c\leq t<c+c^{\prime}},\mathcal{E})\\ &\cong\prod_{c\in\mathbb{R}/\mathbb{G}}\operatorname{Hom}_{\mathrm{Sh}(\mathbb{R}_{t})}(\mathbb{K}_{c+\widetilde{c}\leq t<c+\widetilde{c}+c^{\prime}},\mathcal{E})\end{split}$ (5.18)
where $\widetilde{c}$} is a representative of $c\in\mathbb{R}/\mathbb{G}$. Hence the microstalk of $\mathcal{E}$ at any positive codirections vanish. Hence $\mathcal{E}$ has non-positive microsupport. This implies $\mathcal{E}=0$, since $\mathcal{E}$ is an object of (equivariant) Tamarkin category.
@@
$\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\operatorname{Hom}(\mathbb{K}_{(-\infty,0)},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow.$ (5.22)
As in the proof of Lemma 5.2, we can see that $H^{1}(\operatorname{Hom}(\mathbb{K}_{\mathbb{R}},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c}))\simeq 0$. Then one can go to the cohomology exact sequence as
$0\rightarrow H^{0}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow\widehat{\bigoplus}_{c\in(-\infty,+\infty),I}\mathbb{K}\rightarrow\widehat{\bigoplus}_{c\in(-\infty,0),I}\mathbb{K}\rightarrow H^{1}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\rightarrow 0.$ (5.23)
Here $\widehat{\bigoplus}_{c\in(-\infty,+\infty),I}\mathbb{K}$ is the subspace of $\prod_{c\in(-\infty,+\infty),I}\mathbb{K}$ spanned by the elements of the form $\prod_{c\in(-\infty,+\infty),i\in I}a_{c,i}$ ($a_{c,i}\in\mathbb{K}$) satisfying
$\left\{(c,i)\in(-\infty,N)\times I\mathrel{}\middle|\mathrel{}a_{c,i}\neq 0\right\}<+\infty$ (5.24)
for any $N\in\mathbb{R}$. Similarly, $\widehat{\bigoplus}_{c\in(-\infty,0),I}\mathbb{K}$ is the subspace of $\prod_{c\in(-\infty,0),I}\mathbb{K}$ spanned by the elements of the form $\prod_{c\in(-\infty,0),i\in I}a_{c,i}$ ($a_{c,i}\in\mathbb{K}$) satisfying
$\left\{(c,i)\in(-\infty,N)\times I\mathrel{}\middle|\mathrel{}a_{c,i}\neq 0\right\}<+\infty$ (5.25)
for any $N\in\mathbb{R}_{<0}$. This computation is done in the exactly same manner as the proof of Lemma 5.1.3, where its proof can be found in [Kuw24].
Hence we conclude that
$H^{i}\operatorname{Hom}(\mathbb{K}_{t\geq 0},\bigoplus_{I}\bigoplus_{c\in\mathbb{R}}\mathbb{K}_{t\geq c})\cong\begin{cases}&\Lambda_{0}\text{ if $i=0$}\\ &\operatorname{coker}\left(\bigoplus_{(-\infty,0),I}\mathbb{K}\rightarrow\widehat{\bigoplus}_{(-\infty,0),I}\mathbb{K}\right)\text{ if $i=1$}\\ &0\text{ otherwise}.\end{cases}$ (5.26)
where the morphism in the second line is a natural one. Then the almostization kills the degree one morphisms.
@@
$\operatorname{Hom}(1_{\mu},\bigoplus_{i\in I}T_{d_{i}}1_{\mu})\cong_{\mathrm{a}}\widehat{\bigoplus}_{i\in I}\operatorname{Hom}(T_{c}1_{\mu},T_{d_{i}}1_{\mu}).$ (5.27)
@@
With Step 1, the argument of [GR17, Proposition 5.4.5] immediately shows that $\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu}$ generates the whole category by colimits. Since $\mathfrak{A}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong_{\mathrm{a}}L_{0}^{\mathbb{G}}$ by Lemma 5.3 and
$\operatorname{End}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu})\cong_{\mathrm{a}}L_{0}^{\mathbb{G}}\cong\operatorname{End}(L_{0}^{\mathbb{G}}),$ (5.28)
we conclude that the functor $\mathfrak{A}$ is almost fully faithful by using Step 2.
Also, since $L_{0}^{\mathbb{G}}$ is a generator of $\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$ and $\mathfrak{A}$ is cocontinuous, the almost essential surjectivity follows. This completes the proof. ∎
Proposition 5.7 (The inverse functor).
The functor $\mathfrak{B}\colon\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})\to\mu^{\mathbb{G}}(*);M\mapsto M\otimes_{L_{0}^{\mathbb{G}}}\left(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu}\right)$ is a left adjoint of $\mathfrak{A}$ and an almost equivalence. Moreover $\mathfrak{A}\circ\mathfrak{B}(M)\cong_{\mathrm{a}}M,\mathfrak{B}\circ\mathfrak{A}(\mathcal{E})\cong_{\mathrm{a}}\mathcal{E}$ for any $M,\mathcal{E}$.
Proof.
For $M,N\in\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$, we have
$\operatorname{Hom}(M,\mathfrak{A}(\mathcal{E}))=\operatorname{Hom}(M,\operatorname{Hom}(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu},\mathcal{E})\cong\operatorname{Hom}(M\otimes_{L_{0}^{\mathbb{G}}}\left(\bigoplus_{c\in\mathbb{R}/\mathbb{G}}T_{c}1_{\mu}\right),\mathcal{E}).$ (5.29)
Hence $\mathfrak{B}$ is a left adjoint of $\mathfrak{A}$.
The assertion $\mathfrak{A}\circ\mathfrak{B}(M)\cong_{\mathrm{a}}M$ follows from the proof of the almost essential surjectivity of $\mathfrak{A}$. We then have
$\operatorname{Hom}(\mathfrak{B}(M),\mathfrak{B}(N))\cong\operatorname{Hom}(M,\mathfrak{A}\circ\mathfrak{B}(N))\cong_{\mathrm{a}}\operatorname{Hom}(M,N),$ (5.30)
which shows the almost fully faithfulness of $\mathfrak{B}$. Since $\mathfrak{A}$ is almost fully faithful, we have
$\operatorname{Hom}(\mathfrak{B}\circ\mathfrak{A}(\mathcal{E}),\mathcal{F})\cong\operatorname{Hom}(\mathfrak{A}(\mathcal{E}),\mathfrak{A}(\mathcal{F}))\cong_{\mathrm{a}}\operatorname{Hom}(\mathcal{E},\mathcal{F})$ (5.31)
This shows $\mathfrak{B}\circ\mathfrak{A}(\mathcal{E})\cong_{\mathrm{a}}\mathcal{E}$. This also show the almost essential surjectivity of $\mathfrak{B}$. This completes the proof. ∎
@@ 6.1 Reminders on the Lurie tensor product
We follow Volpe’s exposition [Vol25]. Let $\mathrm{Mod}_{\mathrm{Mod}(\mathbb{K})}(Cat_{comp})$ be the category of the cocomplete $\mathbb{K}$-linear categories and the morphisms are $\mathbb{K}$-linear cocontinuous functors. For $\mathbb{K}$-linear cocomplete categories $\mathcal{C},\mathcal{D}$, there exists a $\mathbb{K}$-linear cocomplete category $\mathcal{C}\otimes_{L}\mathcal{D}$ with a functor $\mathcal{C}\times\mathcal{D}\rightarrow\mathcal{C}\otimes_{L}\mathcal{D}$ satisfying
$Fun_{!\times!}(\mathcal{C}\times\mathcal{D},\mathcal{E})\cong Fun_{!}(\mathcal{C}\otimes_{L}\mathcal{D},\mathcal{E})$ (6.1)
where the left hand side $Fun_{!\times!}$ denotes the $\mathbb{K}$-linear functors preserving variable-wise colimits and the right hand side $Fun_{!}$ denotes the $\mathbb{K}$-linear functors preserving colimits. The resulting category $\mathcal{C}\otimes_{L}\mathcal{D}$ is called Lurie’s tensor product. In the following, we simply denote $\otimes_{L}$ by $\otimes$.
@@
Lemma 6.1 ([Vol25, Corollary 2.30]).
Let $X$ and $Y$ be manifolds.
@@
    We have an equivalence. $\mathrm{Sh}(X,\mathbb{K})\otimes\mathrm{Sh}(Y,\mathbb{K})\cong\mathrm{Sh}(X\times Y,\mathbb{K})$.
@@
    Let $\mathcal{C}$ be a presentable $\mathbb{K}$-linear category. Then we have $\mathrm{Sh}(X,\mathbb{K})\otimes\mathcal{C}\cong\mathrm{Sh}(X,\mathcal{C})$.